OpenStudy (anonymous):

Determine whether the graph of y = x2 + 4x − 7 has a maximum or minimum point, then find the maximum or minimum value. Minimum; (-11, -2) Maximum; (-11, -2) Maximum; (-2, -11) Minimum; (-2, -11)

OpenStudy (anonymous):

minimum value

OpenStudy (anonymous):

\[-\frac{b}{2a}=-\frac{4}{2\times 1}=-2\] \[(-2)^2+4\times (-2)-7=4-8-7=-11\]

OpenStudy (anonymous):

i dont knowwwwwww